OpenStudy (anonymous):

(1/2)|5x+2|-7 is less than or equal to -3+2x

OpenStudy (anonymous):

0.5|5x+2|-7≤-3+2x treat as an equation and find both x. positive case: 0.5(5x+2)-7=-3+2x 2.5x+1-7=-3+2x 0.5x=3 /*2 x=6 negative case: 0.5(-(5x+2))-7=-3+2x 0.5(-5x-2)-7=-3+2x -2.5x-1-7=-3+2x 4.5x=-5 /: 4.5 x=-10/9 check if both answers are indeed correct (they are)

OpenStudy (anonymous):

Place the x values on the numbers axis, and find what happens in between the answers by plugging arbitrary numbers in the original inequality.

OpenStudy (anonymous):